In 2018, Laureen is currently single. She paid $2,720 of qualified tuition and related expenses for each of her twin daughters Sheri and Meri to attend State University as freshmen ($2,720 each for a total of $5,440). Sheri and Meri qualify as Laureen’s dependents. Laureen also paid $1,910 for her son Ryan’s (also Laureen’s dependent) tuition and related expenses to attend his junior year at State University. Finally, Laureen paid $1,410 for herself to attend seminars at a community college to help her improve her job skills.

What is the maximum amount of education credits Laureen can claim for these expenditures in each of the following alternative scenarios?

A) Laureen's AGI is $45,000

American opportunity credit XXXX
Lifetime Learning Credit XXXX

B) Laureen's AGI is $45,000 and Laureen paid $12,420 (not $19,10) for Ryan to attend graduate school ( his 5th year, not his junior year)

American opportunity Credit XXXX
Lifetime Learning Credit XXXX
Answer #1

a) Laureen's AGI - $45,000

For 2 daughter - AOTC is - (2000*2child) + (720*25%+2child)

=4000+360

=4360

For Ryan - 1910

AOTC - 6270

Laureen lifetime learning credit - Eligible is 2000 (The amount of the credit is 20 percent of the first $10,000 of qualified education expenses or a maximum of $2,000 per return)

So in the above case, it is - 1410*20% =282 (Since AGI is below clip of 56000 he can claim same)

=6310+282 = 6592 is eligible deduction

b)

For Daughter, it is the same as an above i.e. 4,360

For Ryan it is = 2000+ (10000*25%) or maximum 4000

=2000+2500 or 4000

So 4000 is allowed

So AOTC total of 8360 and LLC of 282 so claimed is 8642